# Do Large Language Models Mirror Cognitive Language Processing?

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable abilities in text comprehension and logical reasoning, indicating that the text representations learned by LLMs can facilitate their language processing capabilities. In cognitive science, brain cognitive processing signals are typically utilized to study human language processing. Therefore, it is natural to ask how well the text embeddings from LLMs align with the brain cognitive processing signals, and how training strategies affect the LLM-brain alignment? In this paper, we employ Representational Similarity Analysis (RSA) to measure the alignment between 23 mainstream LLMs and fMRI signals of the brain to evaluate how effectively LLMs simulate cognitive language processing. We empirically investigate the impact of various factors (e.g., pre-training data size, model scaling, alignment training, and prompts) on such LLM-brain alignment. Experimental results indicate that pre-training data size and model scaling are positively correlated with LLM-brain similarity, [^1] For notational simplicity, we refer to the similarity between LLM representations and brain cognitive language processing signals as LLM-brain similarity. and alignment training can significantly improve LLM-brain similarity. Explicit prompts contribute to the consistency of LLMs with brain cognitive language processing, while nonsensical noisy prompts may attenuate such alignment. Additionally, the performance of a wide range of LLM evaluations (e.g., MMLU, Chatbot Arena) is highly correlated with the LLM-brain similarity.
@@ 1 Introduction
Large language models, e.g., chatGPT (OpenAI,, 2023), have demonstrated linguistic capabilities that progressively approach human-level language comprehension and generation. The text representations learned by LLMs transform text into a high-dimensional semantic space (Mikolov et al.,, 2013). In neuroscience, brain cognitive processing signals, e.g., EEG, fMRI signals, have been widely acknowledged to be able to represent a snapshot of cognitive representations of language in human brains (Xu et al.,, 2016). Such signals record human brain activities during the cognitive processing of language and play a key role in revealing the internal workings of human language cognition (Friederici,, 2011).
@@
Figure 1: Diagram of the proposed LLM-brain similarity estimation framework. ‘S’ denotes sentence representations from different modalities. ‘$\rho$’ denotes the pearson correlation coefficient.
Thus, it is natural to ask: how “similar” are these two different modalities of language representations (LLMs representations and cognitive language processing signals in the brain) ? or do LLMs mirror human cognitive language processing? Answering these questions might facilitate us to interpret LLMs from the perspective of cognitive science.
@@
Previous studies have explored the assessment of consistency between representations learned by pre-trained language models and cognitive processing signals (Anderson et al.,, 2013, 2015; Xu et al.,, 2016; Hollenstein et al., 2019a, ; Abnar et al.,, 2019). Along this research line, recent studies compare human neuroimaging data with representations learned 43 pre-trained language models that are smaller than 1B-sized LLMs, revealing that language models with stronger predictive abilities for the next word exhibit greater alignment with the human brain (Schrimpf et al.,, 2021). However, current LLMs have scaled up their model size by hundreds of times (e.g. LLaMa3-400B+ (Meta,, 2024)), and been strengthened by alignment training methods such as supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) (Ouyang et al.,, 2022). These factors contribute significantly to the enhanced capabilities of LLMs (Wei et al.,, 2022; Zheng et al., 2023b, ). Nevertheless, it is yet to be tested whether the latest LLMs with scales of 7B and above maintain consistency with human cognitive data. Additionally, the impact of pre-training data size, alignment training, prompts, and other training strategies on LLM-brain similarity has not been investigated.
@@
This paper aims to assess the extent to which LLMs mirror cognitive language processing activity and subsequently analyze the impacts of different factors (e.g., pre-training data size, model scaling, alignment training) on such LLM-brain alignment. Specifically, we adopt functional magnetic resonance imaging (fMRI) signals as human cognitive language processing signals, and RSA (Kriegeskorte et al.,, 2008), widely used in neuroscience (Connolly et al.,, 2012; Diedrichsen and Kriegeskorte,, 2017), to compute similarity scores between LLM representations and human cognitive language processing signals. The framework for LLM-brain similarity computation is illustrated in Figure 1.
@@
Our experiments comprise 23 open-source LLMs, trained with various pre-training data size, scaling and training strategies (i.e., pre-training, SFT, RLHF), to investigate the impact of these factors on LLM-brain similarity. We further analyze the impact of prompts on the human intention understanding abilities of LLMs and the consistency between LLM sentiment and human sentiment. Additionally, we investigate the relationship between the capability of LLM evaluations and LLM-brain similarity.
@@
  • We employ human cognitive language processing signals to assess the resemblance between LLMs and human cognition. By analyzing the similarity of 23 LLMs to human cognitive language processing, we investigate the impact of various factors, such as pre-training data size, model scaling, alignment training and prompts, on the LLM-brain similarity.
@@
  • Our findings indicate that LLMs exhibit high similarity to brain cognitive language processing in positive sentiment, suggesting that LLMs may encode more positive sentimental text during training.
@@ 2 Related Work
Brain-Inspired NLP
Human brain contains intricate neural networks and hierarchical structures, that encompass a broader cognitive complexity and generalization, providing a more comprehensive and profound perspective to understand the performance of LLMs in language cognition (Denève et al.,, 2017; Thiebaut de Schotten and Forkel,, 2022). Thus, aligning brain cognitive processing signals with text embeddings to evaluate LLMs has remained a prominent area of research. Numerous studies explore how embedding representations of words and sentences extracted from NLP models correspond to fMRI or MEG recordings (Mitchell et al.,, 2008; Huth et al.,, 2016; Pereira et al.,, 2018). Hollenstein et al., 2019b propose a framework for cognitive word embedding evaluation called CogniVal. This framework assesses the extent to which language model representations reflect semantic information in the human brain across three modalities: eye-tracking, EEG, and fMRI. Xu et al., (2016) introduce a lightweight tool named BrainBench, designed to evaluate the word semantics in distributional models, using fMRI images corresponding to 60 concrete nouns. Toneva and Wehbe, (2019) interpret the differences among 4 language models concerning layer depth, context length, and attention types using fMRI data. By leveraging insights from attention experiments, the performance of syntax-related tasks is enhanced.
@@
Leveraging the superior semantic expressiveness of text embeddings from pre-trained models, a few studies utilize the computational mapping between brain activity and text embeddings to decode human brains (Mitchell et al.,, 2008; Pereira et al.,, 2018; Nishida and Nishimoto,, 2018). Sun et al., (2019) have explored the effectiveness of different types of distributed representation models in sentence-level brain decoding based on fMRI, and found that supervised structured models exhibited remarkable performance in brain decoding. Additionally, Wehbe et al., (2014) drew an analogy between the recurrent neural network language model and the working mechanisms of the brain during discourse comprehension. They utilize the predictive performance of magnetoencephalography on hidden states at different moments to reveal what information is encoded by brain regions at specific times.
@@ 3 Methodology
To measure the extent to which LLMs align with brain cognitive language processing, we employ the RSA (Kriegeskorte et al.,, 2008) method to assess the similarity in representations between human brains and LLMs when presented with the same stimulus text. Furthermore, we investigate how LLM-brain similarity varies with different prompts and analyze resemblances between LLMs and brain cognitive language processing regarding sentimental polarity using crafted prompts.
@@ 3.1 LLM-Brain Similarity Estimation
RSA is a widely-used technique for comparing cross-modal representation spaces and has found extensive applications in cognitive neuroscience (Connolly et al.,, 2012; Diedrichsen and Kriegeskorte,, 2017). In NLP, this method is utilized to calculate the consistency between two different representation spaces encoding information (Abdou et al.,, 2019; Chrupała and Alishahi,, 2019). Moreover, it can be employed to quantify the relationship between neuroscience and neural network models (King et al.,, 2019).
@@
We hence use RSA to measure the LLM-brain similarity. Firstly, we compute the representation dissimilarity matrix (RDM) separately for brain cognitive processing signals and LLMs embeddings. For each pair of stimulus sentences, we calculate the 1 - Pearson correlation coefficient (${\rho}$) of their representations (either LLM representations or brain cognitive processing signals) as dissimilarity, following the previous works (Dwivedi and Roig,, 2019; Luo et al.,, 2022). Consequently, both the human brain (RDMh) and the LLM (RDMm) yield a diagonal symmetric dissimilarity matrix of size ${n*n}$, where ${n}$ is the number of stimulus sentences. We formulate it as follows:
${\rm RDM}=\left[\begin{array}[]{ccc}0&\cdots&1-\rho(S_{1},S_{n})\\ \vdots&\ddots&\vdots\\ 1-\rho(S_{n},S_{1})&\cdots&0\\ \end{array}\right]$ (1)
@@
We then estimate the similarity between RDMm and RDMh using the Pearson correlation coefficient. A high correlation score implies that representations learned by the LLM are highly aligned with brain cognitive processing signals.
${\rm Sim}=\rho({\rm RDM}_{h},{\rm RDM}_{m})$ (2)
3.2 Prompts
@@
Instruction tuning significantly enhances the in-context learning ability of LLMs by understanding task requirements well through task definitions and corresponding examples (Ouyang et al.,, 2022; Peng et al.,, 2023). We aim to verify the effect of prompts on human intentions by examining whether appended prompts increase LLM-brain similarity.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable abilities in text comprehension and logical reasoning, indicating that the text representations learned by LLMs can facilitate their language processing capabilities. In neuroscience, brain cognitive processing signals are typically utilized to study human language processing. Therefore, it is natural to ask how well the text embeddings from LLMs align with the brain cognitive processing signals, and how training strategies affect the LLM-brain alignment? In this paper, we employ Representational Similarity Analysis (RSA) to measure the alignment between 23 mainstream LLMs and fMRI signals of the brain to evaluate how effectively LLMs simulate cognitive language processing. We empirically investigate the impact of various factors (e.g., pre-training data size, model scaling, alignment training, and prompts) on such LLM-brain alignment. Experimental results indicate that pre-training data size and model scaling are positively correlated with LLM-brain similarity, [^1] For notational simplicity, we refer to the similarity between LLM representations and brain cognitive language processing signals as LLM-brain similarity. and alignment training can significantly improve LLM-brain similarity. Explicit prompts contribute to the consistency of LLMs with brain cognitive language processing, while nonsensical noisy prompts may attenuate such alignment. Additionally, the performance of a wide range of LLM evaluations (e.g., MMLU, Chatbot Arena) is highly correlated with the LLM-brain similarity.
Do Large Language Models Mirror Cognitive Language Processing?
Yuqi Ren1, Renren Jin1, Tongxuan Zhang2, Deyi Xiong1thanks:  Corresponding author. 1 College of Intelligence and Computing, Tianjin University 2 College of Computer and Information Engineering, Tianjin Normal University {ryq20, rrjin, dyxiong}@tju.edu.cn txzhang@tjnu.edu.cn
@@ 1 Introduction
Large language models, e.g., chatGPT (OpenAI, 2023), have demonstrated linguistic capabilities that progressively approach human-level language comprehension and generation. The text representations learned by LLMs transform text into a high-dimensional semantic space (Mikolov et al., 2013). In neuroscience, brain cognitive processing signals, e.g., EEG, fMRI signals, have been widely acknowledged to be able to represent a snapshot of cognitive representations of language in human brains (Xu et al., 2016). Such signals record human brain activities during the cognitive processing of language and play a key role in revealing the internal workings of human language cognition (Friederici, 2011). Thus, it is natural to ask: how “similar” are these two different modalities of language representations (LLMs representations and cognitive language processing signals in the brain)? Or do LLMs mirror human cognitive language processing? Answering these questions might facilitate us to interpret LLMs from the perspective of neuroscience.
@@
Figure 1: Diagram of the proposed LLM-brain similarity estimation framework. ‘H’ denotes sentence representations from different modalities. ‘$\rho$’ denotes the pearson correlation coefficient. ‘S’ denotes similarity measurement method.
@@
Previous studies have explored the assessment of consistency between representations learned by pre-trained language models and cognitive processing signals (Xu et al., 2016; Hollenstein et al., 2019; Abnar et al., 2019). Along this research line, recent studies compare human neuroimaging data with representations learned by 43 pre-trained language models that are smaller than 1B-sized LLMs, revealing that language models with stronger predictive abilities for the next word exhibit greater alignment with the human brain (Schrimpf et al., 2021). However, current LLMs have scaled up their model size by hundreds of times (e.g. LLaMa3-400B+ (Meta, 2024)), and been strengthened by alignment training methods such as supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022). These factors contribute significantly to the enhanced capabilities of LLMs (Wei et al., 2022; Zheng et al., 2023b). Nevertheless, it is yet to be tested whether the latest LLMs with scales of 7B and above maintain consistency with human cognitive data. Additionally, the impact of pre-training data size, alignment training, prompts, and other training strategies on LLM-brain similarity has not been investigated.
@@
This paper aims to assess the extent to which LLMs mirror cognitive language processing activity and subsequently analyze the impacts of different factors (e.g., pre-training data size, model scaling, alignment training) on such LLM-brain alignment. Specifically, we adopt functional magnetic resonance imaging (fMRI) signals as human cognitive language processing signals, and RSA (Kriegeskorte et al., 2008), widely used in neuroscience (Connolly et al., 2012; Diedrichsen and Kriegeskorte, 2017), to compute similarity scores between LLM representations and human cognitive language processing signals. The framework for LLM-brain similarity computation is illustrated in Figure 1.
@@
Our experiments comprise 23 open-source LLMs, trained with various pre-training data size, scaling and training strategies (i.e., pre-training, SFT, RLHF), to investigate the impact of these factors on LLM-brain similarity. We further analyze the impact of prompt strategy on the human intention understanding abilities of LLMs and the consistency between LLM sentiment and human sentiment. Additionally, we investigate the relationship between the capability of LLM evaluations and LLM-brain similarity.
@@
  • We employ human cognitive language processing signals to assess the resemblance between LLMs and human cognition. By analyzing the similarity of 23 LLMs to human cognitive language processing, we investigate the impact of various factors, such as pre-training data size, model scaling, alignment training and prompt strategy, on the LLM-brain similarity.
@@
  • Our findings indicate that LLMs exhibit high similarity to human cognitive language processing in positive sentiment, suggesting that LLMs may encode more positive sentimental text during training.
@@ 2 Related Work
Neuroscience-Inspired NLP
Human brain contains intricate neural networks and hierarchical structures, that encompass a broader cognitive complexity and generalization, providing a more comprehensive and profound perspective to understand the performance of LLMs in language cognition (Denève et al., 2017; Thiebaut de Schotten and Forkel, 2022; Ren and Xiong, 2021). Thus, aligning brain cognitive processing signals with text embeddings to evaluate LLMs has remained a prominent area of research. Numerous studies explore how embedding representations of words and sentences extracted from NLP models correspond to fMRI or MEG recordings (Mitchell et al., 2008; Pereira et al., 2018). Hollenstein et al. (2019) propose a framework for cognitive word embedding evaluation called CogniVal. This framework assesses the extent to which language model representations reflect semantic information in the human brain across three modalities: eye-tracking, EEG, and fMRI. Xu et al. (2016) introduce a lightweight tool named BrainBench, designed to evaluate the word semantics in distributional models, using fMRI images corresponding to 60 concrete nouns. Toneva and Wehbe (2019) interpret the differences among 4 language models concerning layer depth, context length, and attention types using fMRI data. By leveraging insights from attention experiments, the performance of syntax-related tasks is enhanced.
@@
Leveraging the superior semantic expressiveness of text embeddings from pre-trained models, a few studies utilize the computational mapping between brain activity and text embeddings to decode human brains (Mitchell et al., 2008; Pereira et al., 2018; Ren and Xiong, 2022). Sun et al. (2019) have explored the effectiveness of different types of distributed representation models in sentence-level brain decoding based on fMRI, and found that supervised structured models exhibited remarkable performance in brain decoding. Additionally, Wehbe et al. (2014) drew an analogy between the recurrent neural network language model and the working mechanisms of the brain during discourse comprehension. They utilize the predictive performance of magnetoencephalography on hidden states at different moments to reveal what information is encoded by brain regions at specific times.
@@ 3 Methodology
To measure the extent to which LLMs align with human cognitive language processing, we employ the RSA (Kriegeskorte et al., 2008) method to assess the similarity in representations between human brains and LLMs when presented with the same stimulus text. Furthermore, we investigate how LLM-brain similarity varies with different prompts and analyze resemblances between LLMs and human cognitive language processing regarding sentimental polarity using crafted prompts.
@@ 3.1 LLM-Brain Similarity Estimation
RSA is a widely-used technique for comparing cross-modal representation spaces and has found extensive applications in neuroscience (Connolly et al., 2012; Diedrichsen and Kriegeskorte, 2017). In NLP, this method is utilized to calculate the consistency between two different representation spaces encoding information (Abdou et al., 2019; Chrupała and Alishahi, 2019). Moreover, it can be employed to quantify the relationship between human brains and neural network models (King et al., 2019).
@@
We hence use RSA to measure the LLM-brain similarity. Firstly, we compute the representation dissimilarity matrix (RDM) separately for brain cognitive processing signals and LLMs embeddings. For each pair of stimulus sentences, we calculate the 1 - Pearson correlation coefficient (${\rho}$) of their representations (either LLM representations or brain cognitive processing signals) as dissimilarity, following the previous works (Dwivedi and Roig, 2019; Luo et al., 2022). Consequently, both the human brain (RDMh) and the LLM (RDMm) yield a diagonal symmetric dissimilarity matrix of size ${n*n}$, where ${n}$ is the number of stimulus sentences. We formulate it as follows:
${\rm RDM}=\left[\begin{array}[]{ccc}0&\cdots&1-\rho(H_{1},H_{n})\\ \vdots&\ddots&\vdots\\ 1-\rho(H_{n},H_{1})&\cdots&0\\ \end{array}\right]$ (1)
@@
We then estimate the similarity between RDMm and RDMh using the Pearson correlation coefficient, Euclidean distance, Cosine distance and Spearman correlation coefficient. In addition to Euclidean distance, a high correlation score implies that representations learned by the LLM are highly aligned with brain cognitive processing signals.
${\rm Sim}=S({\rm RDM}_{h},{\rm RDM}_{m})$ (2)
3.2 Prompt Strategy
@@
Instruction tuning significantly enhances the in-context learning ability of LLMs by understanding task requirements well through task definitions and corresponding examples (Ouyang et al., 2022; Peng et al., 2023). We aim to verify the effect of prompts on human intentions by examining whether appended prompts increase LLM-brain similarity.